OpenStudy (anonymous):

Do the ratios form a proportion? Explain why or why not. Show all work. 2/7 and 10/35

OpenStudy (xguardians):

Yes, if you multiply \[\frac{ 2 }{ 7 } * 5\] that is equal to\[\frac{ 10 }{ 35 }\]

OpenStudy (xguardians):

If you multiply the bottom and top is what I meant to say.

OpenStudy (anonymous):

To compare ratios first see the numerator and denominator seperately. numerators are 2 and 10 Which is 2*5 =10 denominators 7 and 5 which is 7*5 =35 If we multiply the numerator and denominator of any fraction by the same number then the resulting fraction is a proportion